A ball is thrown straight up with a speed of 30 m/s.
(a.) How long does it take the ball to reach the maximum height?
(b.) What is the maximum height reached by the ball?
(c.) What is its speed after 4.2 s?
1
Expert's answer
2020-04-06T09:08:09-0400

a)


"t=\frac{v}{g}=\frac{30}{9.8}=3.1\ s"

b)


"h=\frac{v^2}{2g}=\frac{30^2}{2(9.8)}=46\ m"

c) Velocity:


"V=v-gT=30-(9.8)(4.2)=-11\frac{m}{s}"

Negative sign means that a ball moves downwards.

Speed:


"|V|=11\frac{m}{s}"